OpenStudy (anonymous):

Find the vertex of the parabola formed by the following quadratic function: y = -3x2 + 12x – 7 (-2,-43) (2, 5) (-1, -22) (1, 2)

OpenStudy (sburchette):

The x coordinate of the vertex of a parabola is given by\[\frac{-b}{2a}\]In this parabola, the a=-3 and b=12. Substituting we get, \[\frac{-12}{2(-3)}\] this simplifies to -12/-6=2 So the x coordinate is 2. Substituting 2 for x will give the y value. -3(2)^2+12(2)-7. -3(4)+24-7=-7. The vertex is at (2,-7)

OpenStudy (anonymous):

the answer is not up there

OpenStudy (anonymous):

the first coordinate of the vertex is 2, but the second is \[-3(2)^2-12\times 2-7=-12+24-7=12-7=5\]

OpenStudy (sburchette):

Oops, messed up my arithmetic. My apologies.

OpenStudy (anonymous):

typo in my answer \[-3(2)^2+12\times 2-7=-12+24-7=12-7=5\]